### v1

# Every Pixel Tells a Story: End-to-End Urdu Newspaper OCR

Abstract
This paper introduces a comprehensive end-to-end pipeline for Optical Character Recognition (OCR) on Urdu newspapers. In our approach, we address the unique challenges of complex multi-column layouts, low-resolution archival scans, and diverse font styles. Our process decomposes the OCR task into four key modules: (1) article segmentation, (2) image super-resolution, (3) column segmentation, and (4) text recognition. For article segmentation, we fine-tune and evaluate YOLOv11x to identify and separate individual articles from cluttered layouts. Our model achieves a precision of 0.963 and mAP@50 of 0.975. For super-resolution, we fine-tune and benchmark the SwinIR model (reaching 32.71 dB PSNR) to enhance the quality of degraded newspaper scans. To do our column segmentation, we use YOLOv11x to separate columns in text to further enhance performance - this model reaches a precision of 0.970 and mAP@50 of 0.975. In the text recognition stage, we benchmark a range of LLMs from different families, including Gemini, GPT, Llama, and Claude. The lowest WER of 0.133 is achieved by Gemini-2.5-Pro.
Every Pixel Tells a Story: End-to-End Urdu Newspaper OCR
Samee Arif*  and Sualeha Farid* University of Michigan - Ann Arbor asamee@umich.edu, sualeha@umich.edu
11footnotetext: These authors contributed equally to this work.
1 Introduction
OCR plays a critical role in the digitization of printed material, enabling large-scale archival, search, and accessibility applications across languages and domains (Adenekan, 2024). In particular, newspaper OCR is essential for preserving records, supporting digital journalism, and making large volumes of textual data machine-readable. However, OCR for newspapers presents unique challenges due to complex layouts, variable font styles, and poor print quality.
These challenges are further compounded in low-resource languages like Urdu. Urdu script is inherently cursive, with context-sensitive letter forms, frequent ligatures, and baseline variations that make character segmentation and recognition difficult (Zaki et al., 2019). Urdu newspapers introduce an additional layer of complexity with multi-article, multi-column cluttered layouts, and stylized fonts, all of which degrade the performance of conventional OCR pipelines.
This paper introduces an end-to-end pipeline consisting of (1) Image Segmentation (article and column based), (2) Image Super Resolution, and (3) Text Recognition for OCR of Urdu newspapers, addressing the aforementioned challenges. Our contributions can be summarized as follows:
  1. 1.
    We fine-tune two separate YOLOv11x (Redmon et al., 2016) models: one for article-level segmentation and another for column-level segmentation, addressing the multi-column structure of Urdu newspaper layouts.
  2. 2.
    We fine-tune SwinIR (Liang et al., 2021) for image enhancement, improving degraded archival scans via super-resolution.
  3. 3.
    We benchmark LLMs including Gemini-2.5-Pro (Google et al., 2024), GPT-4o and 4.1 (OpenAI et al., 2024), Llama-4-Scout and Maverick (Meta, ), and Claude-3.7-Sonnet (Anthropic, ) for Urdu text recognition.
  4. 4.
    We release the dataset for LLM benchmarking and the first publicly available models for article and column segmentation for newspapers, and super-resolution for Urdu text  [^1] https://github.com/paper-seven/UrduOCR.
2 Related Work
Figure 1: End-to-end OCR pipeline with article and column segmentation, super-resolution, and LLM-based text recognition.
Optical Character Recognition (OCR) for Urdu remains underdeveloped despite its importance for accessibility, digitization, and information retrieval in South Asia. The language poses significant challenges due to its complex cursive script, frequent ligature formations, variable text baselines, and context-dependent glyph shaping. These issues are further compounded by noisy real-world input. As a result, there has been limited success in developing an effective pipeline for Urdu OCR using modern model architectures capable of addressing these challenges.
2.1 Prior Work in Urdu OCR
Early work in Urdu OCR focused on creating segmentation-based approaches such as those by Sabbour and Shafait (2013) who reported an accuracy of 91% on clean Urdu text, and Tabassam et al. (2009) who developed an OCR system for Urdu Naskh font utilizing pattern matching techniques, achieving an accuracy of 89%. However, these methods often struggled with the cursive and context-sensitive nature of the Urdu script.
To address these challenges, segmentation-free approaches were explored. Shabbir and Siddiqi (2016) proposed a system that recognizes Urdu words in the Nastaliq font using ligatures as units of recognition, employing Hidden Markov Models (HMMs) for classification. Similarly, Khattak (2017a) presented a segmentation-free OCR system for printed Urdu text, also using HMMs and focusing on ligature recognition. These approaches demonstrated improved handling of the complex ligature structures inherent in the Urdu script.
2.2 Recent Advances in Multilingual OCR
In recent years, there has been significant advancement in multilingual OCR systems with transformer-based architectures. For example, the MIT-10M dataset by Khattak (2017b) contains 10 million plus image-text pairs in 14 languages. In another work by Wang et al. Wang et al. (2024), the authors proposed the use of LLMs for OCR pipelines and discussed ways to mitigate hallucinations to enhance reliability in multilingual image-to-text pipelines. However, Urdu remains underrepresented in both large-scale datasets and employing current LLM based methods for this task.
3 Methodology
3.1 Datasets
Table 1 summarizes the dataset sizes used for each stage of our pipeline. The OCR evaluation dataset comprises 829 samples and includes 8,785 unique words and 2,539 sentences, demonstrating significant lexical diversity. All transcriptions were generated manually and verified by a second annotator to ensure high accuracy. Unlike traditional OCR datasets that focus on sentence-level inputs, our samples consist of full paragraphs of Urdu text to enable rigorous evaluation of LLMs for text recognition. For the image super-resolution dataset, we generated low-quality inputs by downscaling high-resolution Urdu newspaper scans by a factor of 4 and reducing the quality by 30% using JPEG compression.
Stage Train Size Test Size
Article Segmentation 26,830 2,975
Column Segmentation 3969 440
Super-Resolution 38,512 4,279
Text Recognition - 829
Table 1: Dataset sizes used for article segmentation, image super-resolution, and text recognition.
3.2 Pipeline
Our end-to-end OCR pipeline consists of four main stages: article segmentation, image super-resolution, column segmentation, and text recognition using LLMs. Figure 1 illustrates the complete workflow.
In the first stage, we use YOLOv11x for article-level segmentation. Following this, we apply our super-resolution model to the cropped article images in stage two. By enhancing image clarity and sharpening text details, this step produces cleaner, more uniform text blocks, which in turn enables more accurate column segmentation and more accurate text recognition. The low-quality cropped articles are passed through SwinIR model fine-tuned to increase the image resolution by 4 times. In the third stage we introduce column segmentation following the observation of improved model performance on shorter, single-column text blocks compared to full multi-column articles. This additional step breaks down complex layouts into more manageable units, improving recognition accuracy. In the fourth and final stage, we used LLMs to extract the text from each high-quality image segment.
3.3 Experimental Setup
In this paper, we fine-tune and evaluate YOLOv11x model for article and column level segmentation using standard object detection metrics including precision, recall, mAP@50, and mAP@50:95. The SwinIR model is trained and then evaluated using Peak Signal-to-Noise Ratio (PSNR) to measure reconstruction quality. To quantify the impact of super-resolution, we run our LLMs on both low- and high-resolution images generated by our model, comparing the word error rate (WER) and the character error rate (CER) between the two. For the final OCR stage, we assess LLM performance using WER and CER against human-annotated transcriptions. The system prompt provided to the LLMs is included below.
You are an OCR system. Your job is to transcribe image text exactly as shown, without interpretation, paraphrasing, translation, summarization, or hallucination.
4 Results and Discussion
4.1 Image Segmentation
We evaluate article and column-level segmentation performance using standard object detection metrics as shown in Table 2.
Segmentation
Metric Article Column
Precision 0.963 0.970
Recall 0.971 0.997
mAP@50 0.975 0.975
mAP@50:95 0.866 0.854
Table 2: Scores for article and column segmentation.
Article Segmentation.
For article-level segmentation, our fine-tuned YOLOv11x delivers both high accuracy and coverage. It achieves a precision of 0.963—meaning over 96% of its detected regions are true articles—and a recall of 0.971, capturing more than 97% of all actual article blocks. It also scores an mAP@50 of 0.975 and an mAP@50:95 of 0.866, demonstrating robust detection under both lenient and strict IoU thresholds. Together, these metrics confirm that our model reliably isolates article regions from the complex, multi-column layouts of Urdu newspapers, providing a solid foundation for super-resolution and column segmentation.
Column Segmentation.
Our fine-tuned YOLOv11x demonstrates outstanding performance on column-level segmentation. The precision of 0.970 indicates that 97.0% of all predicted column regions correspond to actual columns, minimizing false positives. The recall of 0.997 indicates that 99.7% of true column regions are successfully detected, ensuring almost no columns are missed. Under a loose IoU threshold (mAP@50), the model achieves 0.975, reflecting high tolerance for localization variation. Even when evaluated under the stricter mAP@50:95 metric, it maintains a robust score of 0.854, showing consistent accuracy across tighter overlap requirements. These results confirm that the model can reliably identify and localize column segments in complex Urdu newspaper layouts, providing a solid foundation for subsequent text extraction tasks.
4.2 Image Super-Resolution
After 100 epochs of training article-segmented data, SwinIR achieves a 32.71 dB PSNR on the test set. This level of image reconstruction is noteworthy given the challenges posed by Urdu script. A PSNR above 30 dB is generally considered strong for image restoration tasks (Kim et al., 2025; Anandhi and Mahalingam, 2025). In the context of Urdu OCR, it shows that the textual feature such as stroke continuity and glyph boundaries are effectively preserved which is essential for accurate downstream text recognition. Appendix A presents visual examples of low- and high-resolution Urdu newspaper images, demonstrating the improvements achieved by our model.
As shown in Table 3, our super-resolution model significantly improves text recognition performance across LLMs. Notably, Gemini-2.5-Pro, which already performs well on low-resolution inputs with a WER of 0.177, still benefits from enhancement, reducing its WER further to 0.133. Both GPT-4.1 and GPT-4o also show marked improvements, with WERs dropping from 0.682 to 0.254 and 0.779 to 0.327, respectively. Llama-4-Maverick also exhibit dramatic gains, decreasing WER from 1.036 to 0.305 after super-resolution. Claude-3.7-Sonnet and Llama-4-Scout fail to process most of the low-resolution images, often returning messages such as "Unfortunately, I am unable to extract text from the image…", "I can’t directly extract text…when the image quality does not allow for clear text recognition.", and "The image contains text in what appears to be Urdu script…" instead of performing any transcription.
4.3 Text Recognition
We evaluate text recognition performance of LLMs on the super-resolved images using WER and CER, as reported in Table 3.
Low-Res High-Res
LLM WER CER WER CER
GPT-4.1 0.682 0.471 0.254 0.096
GPT-4o 0.779 0.559 0.327 0.154
Gemini-2.5-Pro 0.177 0.046 0.133 0.032
Claude-3.7-Sonnet Fail Fail 0.249 0.100
Llama-4-Maverick 1.036 0.837 0.305 0.128
Llama-4-Scout Fail Fail 0.347 0.164
Table 3: WER before and after passing images through super-resolution model.
Each model was prompted using the following user message to ensure accurate text recognition without any unwanted additions:
Extract the exact text from this image. Preserve sentence structure NOT spacing. If anything is unreadable, write ’[UNREADABLE]’.
For high-resolution images, Gemini-2.5-Pro outperforms all other models, achieving the lowest WER and CER of 0.133 and 0.032, respectively. This indicates strong performance not only at the word level but also in character-level precision, making it well-suited for accurately transcribing Urdu. GPT-4.1 ranks as a close second, with a WER of 0.254 and CER of 0.096. While Claude-3.7-Sonnet struggles in low-resolution settings, it performs well on high-resolution images, achieving a WER of 0.249 and CER of 0.100, suggesting that its performance benefits significantly from our super-resolution module.
In contrast, GPT-4o underperforms with a WER of 0.327 and CER of 0.154, possibly due to overfitting to English structured text formats and reduced robustness for Urdu. Both Llama-4 models show relatively weaker performance, with Maverick outperforming Scout. Maverick records a WER of 0.305 and CER of 0.128, while Scout lags behind with a WER of 0.347 and CER of 0.164, aligning with expectations given Maverick’s larger model size.
5 Conclusion
Our proposed pipeline enables a mess-to-text transformation: complex, low-quality Urdu newspaper images are converted into clean, human-readable text without any manual correction or user intervention. By integrating fine-tuned models for article and column segmentation with super-resolution enhancements, we significantly improve the quality of inputs and therefore the results of large language models for OCR tasks.
These improvements lead to substantial performance gains in text recognition, significantly reducing WER and CER across models. Notably, they enable models like Claude-3.7-Sonnet, which previously failed under low-resolution conditions, to complete the OCR task successfully. This highlights the importance of layout-aware preprocessing and image enhancement for effective LLM-based approaches in low-resource settings.
Overall, we present a robust, modular system for high-quality Urdu OCR that significantly enhances the performance of general-purpose LLMs on complex scripts. Our pipeline demonstrates how targeted preprocessing can unlock capabilities of general purpose models for low-resource languages like Urdu in such tasks. We also release a curated text recognition evaluation dataset to support further research in this area.
6 Future Work
Moving forward, we plan to (i) expand our Urdu text recognition corpus and fine-tune open-source LLMs such as Gemma-3-12B and Gemma-3-27B on this data; (ii) develop OCR post-processing pipelines using LLMs to automatically correct residual errors and normalize transcriptions; and (iii) extend our evaluation of LLM-based text recognition to additional regional scripts—specifically Punjabi, Sindhi, and Balochi—to assess cross-language generalization and identify language-specific adaptation needs.
Limitations
Our evaluation dataset is relatively small, only 829 manually transcribed Urdu newspaper samples, which limits the statistical power and diversity of our OCR benchmarks (human annotation makes scaling difficult). We rely on off-the-shelf LLMs in a zero-shot setting; fine-tuning these models on in-domain OCR data could yield substantial gains but requires additional annotated text. Finally, the end-to-end pipeline incurs significant computational overhead (multiple YOLOv11x models, SwinIR training, and LLM inference), which may challenge deployment in resource-constrained environments.
Ethical Considerations
While our pipeline advances digitization of Urdu newspapers, several ethical aspects merit attention. First, the source material may contain copyrighted or sensitive content; users must ensure publication rights or fair-use compliance before processing archival or proprietary texts. Second, OCR errors risk propagating misinformation if left uncorrected. Finally, computational demands of large-scale segmentation, super-resolution, and LLM inference can exacerbate environmental impact; practitioners should balance model complexity with energy considerations and, where possible, leverage efficient hardware or model distillation.
Appendix A Super-Resolution Results
Figure 2 shows qualitative results from our super-resolution model. Each row displays a pair of input and output images of our super-resolution model.
(a) Input
(b) Output
(c) Input
(d) Output
Figure 2: Side-by-side comparisons of input and output of super-resolution model.
Appendix B Computing Infrastructure
We use the API for GPT-4o and GPT-4o-mini from OpenAI [^2] https://platform.openai.com/docs/overview. For Gemma and Llama models API from DeepInfra [^3] https://deepinfra.com/ was used. For Claude-3.7-Sonnet, we used their own API [^4] https://www.anthropic.com/api. Finally for Gemini-2.5-Pro, we use Google’s API [^5] https://ai.google.dev/. For YOLO and SwinIR fine-tuning, we use Tesla V100 on VastAI [^6] https://vast.ai/.
Appendix C Human Transcribers
The human transcribers who transcribed the datasets are native Urdu speakers. There were two annotators, and both worked collaboratively, revising the transcriptions together to ensure accuracy and consistency throughout the dataset.
Appendix D Models
YOLOv11 is released under the AGPL-3.0 license, SwinIR is available under Apache-2.0 license, Llama-4-Scout and Llama-4-Maverick are distributed under the LLaMA 4 license. GPT-4o and GPT-4.1 are offered under OpenAI’s proprietary license, while Gemini and Claude are under Google’s and Anthropic’s proprietary licenses, respectively. All of these models allow use for research purposes under their respective licenses.